1 2by3 litres of milk cost rupees 30, what is the cost of 3 5by9 liters of milk?
1
Expert's answer
2019-10-09T09:58:31-0400

If 1 2by3 litres of milk cost rupees 30 then 1 litre costs 30 / (1 + 2/3) = 30/(5/3) = 3*30/5 = 3*6 = 18 rupees

Then 3 5by9 liters of milk will cost 18*(3 + 5/9) = 18*(32/9) = 2*32 = 64 rupees


Answer: 64 rupees